OpenStudy (mikithemermaid):

16y^2 = (xy)² plz help need in the next few minutes thanks

OpenStudy (tkhunny):

You'll have to suggest what is wanted and what you have done to solve it.

OpenStudy (mikithemermaid):

Sorry I need to find the value of x :)

OpenStudy (tkhunny):

Well, give it a go. Looks like y^2 is wild. Get that out of there.

OpenStudy (mikithemermaid):

16=x^2?

OpenStudy (mikithemermaid):

If 16=x^2 would x=8?

rishavraj (rishavraj):

4 * 4 = 16

OpenStudy (mikithemermaid):

Oh so x=4?

OpenStudy (mikithemermaid):

@rishavraj

rishavraj (rishavraj):

yup

OpenStudy (mikithemermaid):

Kay thanks!!!

OpenStudy (tkhunny):

Where did -4 go?